write the 6 trig values in exact form for the point: p(3,4)
Answer by fractalier(6550)

Call the angle the point p makes with the origin q.
Then the trinagle that is formed is a 3-4-5 right triangle and we have
sin q = 4/5
cos q = 3/5
tan q = 4/3
csc q = 5/4
sec q = 5/3
cot q = 3/4